OpenStudy (anonymous):

MEDAL!! (refer to choosing two cards from a thoroughly shuffled deck. Assume that the deck is shuffled after a card is returned to the deck) If you put the first card back in the deck before you draw the next, what is the probability that the first card is a 10 and the second card is a jack?

jimthompson5910 (jim_thompson5910):

Focus on one card at a time. What is the probability of pulling out a "10" card ?

jimthompson5910 (jim_thompson5910):

here's a visual of all the cards (52 total; 4 suits, 13 per suit) http://www.jfitz.com/cards/classic-playing-cards.png

OpenStudy (anonymous):

hmm, there are four 10s

jimthompson5910 (jim_thompson5910):

yes, so the probability of picking a single 10 is?

OpenStudy (anonymous):

guessing 4 out of 52

jimthompson5910 (jim_thompson5910):

you can reduce that fraction to get ?

jimthompson5910 (jim_thompson5910):

4/52 = ???

OpenStudy (anonymous):

1/13

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

so if you focus on one suit only, there is a 1 in 13 chance to get a "10" card

jimthompson5910 (jim_thompson5910):

next step: find the probability of picking a jack. All 52 cards are still there since the first card was put back (it says `Assume that the deck is shuffled after a card is returned to the deck`)

OpenStudy (anonymous):

if there are 4 jacks wouldn't it be 4/52=1/13 again?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

now multiply the two fractions

jimthompson5910 (jim_thompson5910):

P(10 on first draw, jack on second draw) = P(10 on first draw) * P(jack on second draw) this works because the two events are independent

OpenStudy (anonymous):

multiplying the two fractions would get me 1/169

jimthompson5910 (jim_thompson5910):

which is the correct final answer

OpenStudy (anonymous):

thanks again jim!